In a USA Today survey, 510 people chose to respond to the following question that was posted on the USA Today website: “Should Americans replace passwords with biometric security (fingerprints, etc.)?” Among the respondents, 53% said “yes.” We want to test the claim that more than half of the population believes that passwords should be replaced with biometric security.


1.) Are any of the requirements for testing a claim about a population proportion violated? Can the methods of this section be used to test the claim?
Answer #1

1. For using Normal distribution

n (sample size) should be large (n=510 )

and both np and n(1-p) should be greater than 5 (np = 510*53 = 270.3 , n(1-p) = 239.7)

So it fulfill the requirements to use Normal distribution for testing proportions

The null and alternative hypothesis

H0 : p=0.50

Ha: p > 0.50

test statistic,

  

= 1.36

P value =0.0869 (one tailed)

Since P value > 0.05

The result is not significant

There is not enough evidence to conclude that more than half of the population believe that password should be replaced with biometric security